a chemistry student obtained a yield of 3.80 g of H2O(l), at room temperature, from the reaction of 1.60 g H2(g) with 10.0 g O2 (g). what was her percentage yield of this reaction? (Hint: write a balanced reaction equation first and determine the limiting reagent?
1
Expert's answer
2020-09-14T08:17:34-0400

2H2+O2=2H2O

n(H2)=m(H2)/M(H2)=1.6 g/2 g/mol = 0.8 mol

n(O2)=m(O2)/M(O2)=10 g/32 g/mol = 0.3125 mol; for reacting with 0.3125 mol O2 we need only 0.625 mol H2; O2 - limiting reagent.

in theory 0.3125 mol O2 after reaction should give us 0.625 mol of H2O - 0.625 * 18 (M H2O)= 11.25 g. Then yield percentage = 3.80/11.25 *100% = 33.78%